1/1

Lord George Gordon Byron ,

Remembrance

’Tis done!—I saw it in my dreams; No more with Hope the future beams; My days of happiness are few: Chill’d by misfortune’s wintry blast, My dawn of life is overcast; Love Hope, and Joy, alike adieu! Would I could add Remembrance too!

Verkinto de tiu ĉi Angla poemo estas Lord George Gordon Byron (∗1788-01-22 – †1824-04-18).

Arg-1108-2235 (2014-12-26 13:50:48)

Tiu ĉi poemo troveblas en http: // www. mykeep. com/ lordbyron/ remembrance. html .

http://www.poezio.net/version?poem-id=1108&version-id=2235